### v2

# ESDD2: Environment-Aware Speech and Sound Deepfake Detection Challenge Evaluation Plan

Abstract
Audio recorded in real‑world environments often contains a mixture of foreground speech and background environmental sounds. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners. To address this gap, we have proposed CompSpoofV2 dataset [^1] CompSpoofV2: https://xuepingzhang.github.io/CompSpoof-V2-Dataset/ and a separation-enhanced joint learning framework [^2] Baseline: https://github.com/XuepingZhang/ESDD2-Baseline. CompSpoofV2 is a large-scale curated dataset designed for component-level audio anti-spoofing, which contains over 250k audio samples, with a total duration of approximately 283 hours. Based on the CompSpoofV2 and the separation-enhanced joint learning framework, we launch the Environment-Aware Speech and Sound Deepfake Detection Challenge (ESDD2) [^3] ESDD2: https://sites.google.com/view/esdd-challenge/esdd-challenges/esdd-2, focusing on component-level spoofing, where both speech and environmental sounds may be manipulated or synthesized, creating a more challenging and realistic detection scenario. The challenge will be held in conjunction with the IEEE International Conference on Multimedia and Expo 2026 (ICME 2026) [^4] ICME 2026: https://2026.ieeeicme.org/.
I Introduction
In real-world recording scenarios, audio often contains a mixture of foreground speech and background sounds. We define such audio as comprising two components: (i) speech, i.e., linguistically meaningful speech produced by the primary speaker, and (ii) environmental sound, i.e., any non‑speech background or non‑target speech. With rapid advances in text‑to‑speech, voice conversion, and other generation models, either component can now be modified independently. For example, replacing the background while leaving the foreground human speech unchanged, or modifying the speech content while preserving the background. Such component-level manipulations are harder to detect, as the remaining unaltered component can mislead the systems designed for whole deepfake audio, and they often sound more natural to human listeners.
To address this gap, we introduced a component-level spoofing audio dataset CompSpoof [16] in 2025. Then, we extended the dataset to CompSpoofV2, which contains more than 250,000 audio clips (around 283 hours) formed by mixing bona fide and spoofed speech and environmental sound from multiple sources. We also proposed a separation-enhanced joint learning framework to explore to address this gap.
To further promote the development of environment-aware speech and sound deepfake detection, we extend our first ESDD challenge[13] and launch the ICME 2026 Environment-Aware Speech and Sound Deepfake Detection Challenge (ESDD2). While our first ESDD challenge focuses on the detection of forged environmental sounds, ESDD2 advances this direction by addressing component-level spoofing, where both speech and environmental sounds may be manipulated or synthesized, posing a more challenging and realistic detection scenario.
II Dataset
Figure 1: ESDD2 Task illustration. An audio clip is first classified as mixed or original; for mixed audio, speech and environmental sound components are separately evaluated for genuineness, resulting in five target classes.
The CompSpoofV2 dataset contains more than 250,000 audio clips (around 283 hours) formed by mixing bona fide and spoofed speech and environmental sound from multiple sources. Each audio belongs to one of five categories, covering all combinations of bona fide and spoofed speech and environmental sound as shown in Table I. The task of this challenge is to classify a given audio clip into one of the five defined classes, as shown in Figure 1.
TABLE I: Five Categories Considered in CompSpoofV2 dataset.
ID Mixed Speech Environment Class Label Description
0 Bona fide Bona fide original Original audio without any manipulation or mixing
1 Bona fide Bona fide bonafide_bonafide Bona fide speech mixed with bona fide environmental sound from other audio
2 Spoofed Bona fide spoof_bonafide Spoofed speech mixed with bona fide environmental sound
3 Bona fide Spoofed bonafide_spoof Bona fide speech mixed with spoofed environmental sound
4 Spoofed Spoofed spoof_spoof Spoofed speech mixed with spoofed environmental sound
The audio sources for each category are summarized in Table II and Table III.
TABLE II: Audio sources for the training and validation sets.
ID Original Source Speech Source Environmental Sound Source
0 AudioCaps[4], VGGSound[2]
1 CommonVoice[1], LibriTTS[15], english-conversation-corpus[5] AudioCaps, TAUUAS[3], TUTSED[6, 7, 8], UrbanSound[10], VGGSound
2 CommonVoice, LibriTTS EnvSDD[14], VcapAV[12]
3 ASV5[11], MLAAD[9] AudioCaps, TAUUAS, TUTSED, UrbanSound, VGGSound
4 ASV5, MLAAD EnvSDD, VcapAV
TABLE III: Audio sources for the evaluation and test sets.
ID Original Source Speech Source Environmental Sound Source
0 AudioCaps, VGGSound
1 CommonVoice, LibriTTS, english-conversation-corpus AudioCaps, TAUUAS, TUTSED, UrbanSound, VGGSound
2 CommonVoice, LibriTTS EnvSDD, VcapAV, New Generated
3 ASV5, MLAAD, New Generated AudioCaps, TAUUAS, TUTSED, UrbanSound, VGGSound
4 ASV5, MLAAD, New Generated EnvSDD, VcapAV, New Generated
The training and validation sets share the same data sources and class distribution. Similarly, the evaluation and test sets share the same data sources and class distribution. Notably, the evaluation and test sets include newly generated audio samples that are unseen in the training and validation sets.
III Baseline
We propose a separation-enhanced joint learning framework as the baseline for this task. The framework first detects potentially spoofed mixtures, then separates the audio into speech and environmental components, which are then processed by speech- and environment-specific anti-spoofing models. Their outputs are fused and mapped to five-class predictions. The separation and anti-spoofing models are jointly trained to preserve spoofing-relevant cues. More detail can be found in [16]. The baseline results are as IV.
TABLE IV: Performance (EER % and F1-score) of baseline model on the validation, evaluation, and test sets of CompSpoofV2.
Original EER Speech EER Env. EER F1-score
Val. set 0.0031 0.0172 0.3766 0.9462
Eval set 0.0174 0.1993 0.4336 0.6224
Test set 0.0173 0.1978 0.4279 0.6327
IV Evaluation Criteria
System performance will be evaluated using the Overall Macro-F1 score across all five target classes, as shown in  (1). The Macro-F1 score is computed as the arithmetic mean of the per-class F1 scores. This ensures all classes contribute equally to the final evaluation, regardless of sample imbalance.
$\text{Macro-F1}=\frac{1}{5}\sum_{i=1}^{5}\text{F1}_{i},$ (1)
where, the $\text{F1}_{i}$ score is calculated as the harmonic mean of precision $P_{i}$ and recall $R_{i}$ for class $i$:
$\text{F1}_{i}=\frac{2\cdot P_{i}\cdot R_{i}}{P_{i}+R_{i}},\quad P_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}},\quad R_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}.$ (2)
$\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ denote the number of true positives, false positives, and false negatives for class $i$, respectively. A higher Macro-F1 score indicates better overall system performance.
In addition, we introduce three auxiliary metrics based on equal error rate (EER): $\mathrm{EER}_{\text{original}}$, $\mathrm{EER}_{\text{speech}}$ and $\mathrm{EER}_{\text{env}}$. $\mathrm{EER}_{\text{original}}$ is the EER for distinguishes the Original ($i=0$) class from all other four classes ($i=\{1,2,3,4\}$. Similarly, $\mathrm{EER}_{\text{speech}}$ is the EER for detecting spoofed speech component, while $\mathrm{EER}_{\text{env}}$ is the EER for detecting spoofed environmental component. The three EER metrics are included for diagnostic and analysis purposes only; they will not be considered for leaderboard ranking.
V Submission Guideline
We use CodaBench platform for results submission [^5] Results Submission: https://www.codabench.org/competitions/12365/.
Participants are expected to compute and submit a text file. We suppose two submission file formats. One is the two-line format as follows. In the file, each line includes the audio ID and the predicted class ID, which is an integer label from 0 to 4 corresponding to the class ID in Table I.
1 UGBFJ3.wav | 0
2 ABCD12.mp3 | 3
The other is the five-line format as follows. In the file, audio ID, the confidence score for predicting the original class, the confidence score for the speech component, the confidence score for the environmental sound component, and the predicted class ID.
1 UGBFJ3.wav | 1.22 | 0.91 | 0.99 | 0
2 ABCD12.mp3 | -1.54 | 0.89 | -1.89 | 3
The challenge has two phases: the preparation phase and the final phase. During the preparation phase, participants submit evaluation set predictions and the leaderboard reflects evaluation set rankings; and the final evaluation phase, where participants submit test set predictions and the leaderboard determines the final ranking. Each team may submit up to ten results in the final evaluation phase, with the best three used for final ranking to ensure fairness. After the competition, we will release the metadata of the evaluation and test sets to support reproducibility and future research.
Participants may use additional public datasets only with the organizers’ prior approval. Requests must be submitted by email to the organizers (xueping.zhang@dukekunshan.edu.cn) before February 20th, 2026. The organizers will publish the list of approved datasets; any dataset not on this list is not permitted for training.
The ESDD2 challenge is one of the Grand Challenges in ICME 2026. Participants to the Grand Challenge are encouraged to submit short papers (up to 4 pages) that, in case of acceptance, will be published in the conference workshop proceedings.
VI Rules
To ensure fairness, the following rules are established:
  • The evaluation and test sets must not be used for training.
  • Publicly available pre-trained models are allowed but must be disclosed.
  • Participants are encouraged to report model parameters and computing resources.
  • The use of any synthetic audio—including audio generated by text-to-Speech models, voice conversion models, or any other audio generation models—is strictly prohibited.
  • Participants may use additional public datasets only with the organizers’ prior approval. Requests must be submitted before February 20, 2026, and the organizers will publish the list of approved datasets; any datasets not on this list are not permitted for training.
VII Registration Process
The following Google Form will be used by participating teams for registration of their respective teams in the challenge [^6] Registration: https://forms.gle/WnkDkyiXtoVNMZRk7.
VIII Important Dates
The detailed timeline is shown in Table V.
TABLE V: ESDD2 Challenge Timeline.
Registration Opens January 10th, 2026
Training, validation and evaluation sets release; Baseline code release; Preparation phase leaderboard update January 30th, 2026
Additional dataset approval deadline February 20th, 2026
Test set release; Final evaluation phase leaderboard update March 20th, 2026
Registration closes and leaderboard freeze April 25th, 2026
Challenge Paper Submission Deadline April 30th, 2026
More information can be found in the Google Group [^7] Google Group: https://groups.google.com/g/esdd2-challenge
IX Sponsors and Awards
OfSpectrum, Inc. (https://ofspectrum.com/), an AI company based in the U.S., is sponsoring this challenge and will provide 1,000 U.S. dollars prize to the first-place winner. OfSpectrum develops imperceptible watermarking technology, with applications in content provenance, copyright protection and facilitating the safe and legitimate use of AI in voice applications.